OpenStudy (anonymous):

In the following reaction, how many grams of ammonia (NH3) will produce 300 grams of N2? 4NH3 + 6NO → 5N2 + 6H2O The molar mass of ammonia is 17.0337 grams and that of nitrogen is 28.02 grams.

OpenStudy (kirbykirby):

In this reaction, 4 moles NH3 produces 5 Moles of N2 1 mole NH3 = 17.0337 g 1 mole N2 = 28.02 g Convert 300 g ammonia into moles: \[ \frac{1 ~mol ~NH_3}{17.0337~g~ NH_3} =\frac{x~ mol NH_3}{300 ~ g ~NH_3} \implies x=17.612~ mol~NH_3\] Now, you can find the amount of moles that will be produced in the reaction by the ratio of 4 moles NH3 produced for 5 moles of N2: \[ \frac{4~mol ~NH_3}{17.612~mol~NH3}=\frac{5~mol~N_2}{x~ mol~N_2}\] Once you find the amount of moles produced of N2, then covert the moles into grams using the molar mass : 1 mole N2 = 28.02 g

OpenStudy (anonymous):

What was the final answer to this?